A sinking fund is established to raise R50000 in ten years time. The fund must also make provision for withdrawals of R800 every half year starting 6 months from now and finishing 3 and half years from now. The one hundred amd twenty monthly payments that must be made over the ten years period into an account paying 6.3%p.a compounded half yearly are equal to?
1
Expert's answer
2020-05-25T21:15:30-0400


1."50000+3.5\times2\times800=55600"


total amount for the formation of a sinking fund 


2.find monthly payment


"FVA=A\times\frac{(1+r)^n-1}{r}"


"55 600=A\times\frac{(1+0.00525)^{120}-1}{0.00525}"


A=333.78